OpenStudy (anonymous):

If g(x) = 2(x − 4), find the value of x if g(x) = 20.

OpenStudy (anonymous):

We know that: \[ g(x)=2(x-4) \]And: \[ g(x)=20 \]We wish to find such \(x\), so, by transitive property of equality: \[ 20=2(x-4)\implies\\ 10=x-4\implies\\ 14=x \]Et voilá.